OpenStudy (aaronandyson):

The speed of light in diamond is 125000kms^-1.What is the refractive index? @ganeshie8

OpenStudy (aaronandyson):

@Directrix

OpenStudy (shamim):

a mew b=Ca/Cb

OpenStudy (aaronandyson):

ca/cb????????????? what??????????

OpenStudy (shamim):

a mew b=refractive index of the medium b with respect the medium a

OpenStudy (shamim):

Ca=light velocity in a medium a

OpenStudy (shamim):

Cb=light velocity in medium b

OpenStudy (shamim):

a=free space b=diamond

OpenStudy (academicgurusinc):

Would this help? Refractive index of diamond= Speed of light in vacuum or air/ Speed of light in diamond

OpenStudy (aaronandyson):

@Directrix

OpenStudy (shamim):

0 mew b=C0/Cb=3*10^8/125000000=?

OpenStudy (academicgurusinc):

Once you know that concept, you can then apply it: Refractive index of diamond= Speed of light in vacuum or air/ Speed of light in diamond V(air)/V(diamond) = 3x10^8/1.25x10^8 = 2.4 as your final answer. Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (aaronandyson):

"0 mew b=C0/Cb=3*10^8/125000000=?" 125000 it had only 3 zeros?

OpenStudy (aaronandyson):

???????????????????????????????????????????????????????????????????????

OpenStudy (aaronandyson):

@Directrix

OpenStudy (perl):

Do you agree that: Refractive index of diamond= Speed of light in vacuum or air/ Speed of light in diamond

OpenStudy (perl):

It's just a definition. its useful because we can look up the refractive index and then know the speed of light through the medium (here a diamond)

OpenStudy (aaronandyson):

?

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (perl):

basically, light slows down when it goes through a diamond. when you see the 'speed of light' that is the speed of light in a vacuum

OpenStudy (perl):

speed of light in a vacuum: $$\Large c = 2.99792458 \times 10^8 m/s $$ most books round this to $$\Large c = 3.00 \times 10^8 m/s $$

OpenStudy (perl):

ok so far?

OpenStudy (aaronandyson):

yes.

OpenStudy (aaronandyson):

@perl